OpenStudy (anonymous):

how many mins have I lived if I am 30 years old.

OpenStudy (anonymous):

1.57785e7

OpenStudy (anonymous):

30 x 365.25 x 24 x 60

OpenStudy (anonymous):

i think it would depend what year you were born if we take leap years into consideration

OpenStudy (anonymous):

born in 1982

OpenStudy (anonymous):

There are 15,724,800 minutes in 30 years.

OpenStudy (anonymous):

just off for a few minutes :P

hartnn (hartnn):

i would rather count how many lives can i live in a minute....

OpenStudy (anonymous):

Because every year is actually 365 and a quarter days, which is where we derive leap years, with a full extra calendar day.

OpenStudy (anonymous):

Hero can you break down the equation

OpenStudy (anonymous):

L.T. is correct, we make leap for convenience of having to count full days otherwise each year is 365 full days and 6 hours

OpenStudy (anonymous):

30 years x 365.25 days in each year x 24 hours per day x 60 minutes per hour

OpenStudy (anonymous):

@L.T. Thanks...☺

OpenStudy (anonymous):

\[=30 \cancel{years} \times \frac{365.25 \cancel{days}}{\cancel{year}} \times \frac{24 \cancel{hours}}{\cancel{day}} \times \frac{60 minutes}{\cancel{hour}}\]

OpenStudy (anonymous):

im confused cliffsedge

OpenStudy (anonymous):

The units you don't want cancel out, leaving only minutes.